# Sharp Thresholds for Temporal Motifs and Doubling Time in Random Temporal Graphs

### v1

@@ 9.2 Lower Bounds
In this section we present a lower bound to match the upper bound of Lemma˜5.5. However, this time our life will be significantly easier as it suffices to consider only doubling time of sets of containing precisely $\frac{n}{2}$ vertices. As before, we need a bound on the doubling time of any specific sets, and so make use of the following result analogous to Lemma˜9.3.
@@
With that established we may now give the main result of the section, a lower bound on the doubling time of $\mathcal{G}^{O}$. See 5.5 The key notion for the proof of this result is the idea of a “blocker vertex”, i.e. a vertex who is not yet reachable by any member of some large set of vertices. By applying Lemma˜9.5 and Markov’s inequality, followed by a simple counting argument, we are able to show that at least a constant fraction of vertices must be blockers around time $\frac{n\log{n}}{2}$. Then, via drift analysis, we are able to show that at least one blocker vertex will not be included in another edge before time $(1-\epsilon)n\log{n}$. Given the fact the doubling time is lower bounded by the time when the last vertex stops being a blocker, and that a vertex must be included in an edge in order to stop being a blocker, we immediately obtain the bound.
@@
Let $X$ be the number of sets of vertices of size $\frac{n}{2}$ which have a doubling time of at most $\frac{(1-\alpha)n\log{n}}{2(1+\alpha)}$ for some $\alpha>0$ small. We know from Lemma˜9.5 and the linearity of expectation that for some constant $c>0$,
@@
It therefore holds that with probability $\frac{1}{\log{n}}$ there are at least $(1-cn^{(1-\alpha)(\alpha+\log{(1-\alpha}))})\binom{n}{\frac{n}{2}}$ sets with a doubling time of at least $\frac{(1-\alpha)n\log{n}}{2(1+\alpha)}$. Now for a set of size $\frac{n}{2}$ to have a doubling time of at least some given time, requires that some vertex outside of the set is not yet able to be reached by any vertex inside the set. We call such an unreachable vertex a “blocker” vertex for that set. Now any individual blocker vertex can block at most $\binom{n-1}{\frac{n}{2}}=\frac{1}{n}\binom{n}{\frac{n}{2}}$ sets (precisely those that do not contain it). Since we know that with high probability there are $(1-n^{-\beta})\binom{n}{\frac{n}{2}}$ sets which are blocked for some $\beta>0$, for any constant $0<d<1$ it holds that there must be at least $dn$ blocker vertices.
@@
There exists at least one blocker vertex that does not receive an edge between time $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}$ and $(1-\epsilon)n\log{n}$ with probability 1o(1)1-o(1)
Let $B$ be a set of $dn$ vertices chosen arbitrarily at time $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}$ to either be a subset or super-set of the set of blocking vertices at that time. Let $Y_{t}$ be the set of vertices from $B$ that have not yet received an edge between time $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}$ and $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}+t$, and $\mathcal{F}_{t}$ be its corresponding filtration. We have the following two bounds, for $t<2n\log{n}$ and any fixed $\gamma>0$ immediately,
@@
The main claim follows immediately from this, consolidation of constant terms and the observation that if there is at least one blocker vertex that has not received an edge, there must be at least one set that has not yet doubled its size.
@@ 9.3 Tidying Up
In this section, we prove our main theorem on the doubling time in the continuous model. In order to achieve this we will convert our upper and lower bounds on the order discretization of a sample from the continuous model back to the undiscretized version. This is made quite easy by the strong concentration of the number of labels occurring within a given interval, in fact since $\psi$ is degenerate we obtain the ever pleasant binomial distribution. We are able, therefore, to obtain the following simple precursor, showing that with high probability the number of labels appearing in the initial interval is highly concentrate.
@@
For $\alpha,\beta>0$ with $\beta$ small the following holds, for any $\mathcal{G}=(V,E,\lambda)$,
$\mathbb{P}\left[\sum_{e\in E}|\lambda^{\prime}_{[0,\frac{\alpha\log{n}}{r(n-1)}]}(e)|\leq\frac{\alpha(1-\beta)n\log{n}}{2}\right]=o(n^{-1}).$
@@
$\mathbb{P}\left[\sum_{e\in E}|\lambda^{\prime}_{[0,\frac{\alpha\log{n}}{r(n-1)}]}(e)|\geq\frac{\alpha(1+\beta)n\log{n}}{2}\right]=o(n^{-1}).$
The proof of the result itself is then simply an application of Lemma˜9.7 to the bounds implied by Lemma˜5.5 and Lemma˜5.4.
Proof 9.8 (Proof of Theorem˜3.5).
@@
Furthermore, we know from Lemma˜9.7, that for any $\delta>0$ the $(1-\epsilon)n\log{n}$th label falls after $\frac{2(1-\delta)(1-\epsilon)\log{n}}{r(n-1)}$ and that the $(1+\epsilon)n\log{n}$th labels falls before $\frac{2(1+\delta)(1+\epsilon)\log{n}}{r(n-1)}$ with probability 1o(1)1-o(1). Thus taking a union bound over the failure probabilities, and $2\alpha>\delta+\epsilon+\delta\epsilon$, we obtain the claim.
@@ 10.1.1 First Moment
First for the case $\beta>1$. By independence we have that $\mathbb{P}{\left[A_{S,\tau}=1\right]}=\prod_{e\in E}\mathbb{P}{\left[\tau(e)\leq l_{e}\right]}$. By linearity, we obtain,
@@
From Lemma˜8.6 we have,
@@
Since $\sum_{\tau\in\mathcal{T}_{S}}\mathbb{P}{[\tau(e)\leq l_{e}]}=\mathbb{E}[\psi]=r$ is both finite and absolutely convergent, by the repeated application of distributivity over absolutely convergent sequences,,
@@
From Lemma˜8.6 we have,
@@ 10.1.2 Second Moment
Proof 10.2 (Proof of Lemma˜8.21 in the case where $\beta>1$).
@@
Proof 10.3 (Proof of Lemma˜8.21 in the case where $\beta=1$).
@@
$\mathbb{P}{[\mathbf{F}_{(1),T}=0]}\leq\frac{(\frac{\sqrt{2}r}{T})^{2|E_{H}|}n^{2|V_{H}|}(2|V_{H}|)^{4|V_{H}|}\left(1+|E_{H}|(T-1)\right)\sum_{I\subseteq H}n^{-|V_{I}|}\left(\frac{2r}{T}\right)^{|E_{I}|}}{\left(r^{|E_{H}|}T^{1-|E_{H}|}|V_{H}|!\binom{n}{|V_{H}|}\right)^{2}}$
@@
$O(1)\cdot\frac{1+|E_{H}|(T-1)}{T^{2}}\sum_{I\subseteq H}n^{-|V_{I}|}\left(\frac{2r}{T}\right)^{|E_{I}|}$
@@ 10.2 Doubling times
$p=\frac{2}{rn(n-1)-2t}\cdot\left(rs(n-s)-\sum_{e\in\text{Cut}}\lambda^{\prime}_{[0,t)}(e)\right)$
For $n$ sufficiently large, any fixed $S$ and $t\leq\alpha n\log{n}$ this is trivially upper bounded for any $\gamma>0$,
@@
The result then follows from a simple rearrangement, that $r\geq 1$ and that
@@
With that established, we can start working with the random temporal graph itself. We define $A_{t}=|\mathcal{B}_{\mathcal{G}^{O}_{[\alpha n\log{n},\infty)}}\left(S,\alpha n\log{n}+t\right)|$ to be the size of the reachability ball centered at $S$ after $t$ rounds. Further we define $B_{i}=\inf\{t\geq 0:A_{t}\geq s+i\}$. We will take a slightly different view (but equivalent) of the process sampling $\mathcal{G}^{O}$: Each edge begins with $r$ tokens. Then on each round an edge is chosen to become active and discard a token with probability proportional to its remaining tokens, with the process concluding once all tokens have been sampled.
@@
We construct the coupling as follows. If there exists a set $T$ such that $\sum_{e\in T\times V\setminus T}\frac{|\lambda^{\prime}_{[0,\alpha n\log{n}]}(e)|}{r|T|(n-|T|)}>\gamma$ we immediately terminate the coupling, declare it to have failed and sample $X_{s}$ independently of $\mathcal{G}$. If this is not the case, we begin on round $\alpha n\log{n}$ and work round by round. On the $t$th subsequent round,
@@
      $q_{t}=\frac{n(n-1)-2\alpha n\log{n}-2t}{n(n-1)}\cdot\frac{(1-\gamma)rs(n-s)}{rs(n-s)-\sum_{e\in S\times V\setminus S}|\lambda^{\prime}_{[0,\alpha n\log{n}]}(e)|}$
@@
By construction, we have that unless the coupling fails $X_{s}$ samples its first successful trial no earlier than an edge is revealed crossing the cut. However, the coupling only fails if Lemma˜9.1 fails and so we have the claim.
@@
We shall chain together ksk-s variations of the previous coupling and use the good behavior of each one to ensure the good behavior of the next. The coupling is divided into ksk-s phases such that, for $i\in\{0,...,k-s-1\}$, the $i$th phase begins on round $B_{i}$. In the $i$th phase we couple $B_{i+1}-B_{i}$ and $X_{s+i}$, and take
@@
to be the set of edges crossing the cut between the reached and unreached vertices. First if either of the following hold, we immediately terminate the coupling and sample $X_{s}+i$ independently at random:
@@
    There exists a set $T$ such that $\sum_{e\in T\times V\setminus T}\frac{|\lambda^{\prime}_{[0,\alpha n\log{n}+B_{i}]}(e)|}{r|T|(n-|T|)}>\gamma$
@@
    $B_{i}>(\alpha+10)n\log{n}$
If neither of these occur, we work round by round starting at $\alpha n\log{n}+B_{i}$ and on the $t$th subsequent round.
@@
      $q_{t}=\frac{n(n-1)-2\alpha n\log{n}-2t-2B_{i}}{n(n-1)}\cdot\frac{(1-\gamma)rs(n-s)}{rs(n-s)-\sum_{e\in S\times V\setminus S}|\lambda^{\prime}_{[0,\alpha n\log{n}+B_{i}]}(e)|}$
@@
  • If an edge has been revealed from $\text{Cross}(i)$, we sample the $t$th trial for $X_{s+i}$ according to its standard success probability, indepently of $\mathcal{G}$.
Now note that unless the coupling fails for some $i\in\{0,...,k-s-1\}$ we have that $B_{i+1}-B_{i}<X_{s+i}$ for all of the same $i$, and thus $B_{k-s}<\mathbf{\bar{X}}$. It therefore remains to show that the coupling does not fail with high probability. We begin by observing that it holds that with probability $1-o(n^{-1})$, we have that $\mathbf{\bar{X}}<\mathbf{X}\leq 10n\log{n}$ (previous claim) and that no cut has had more than a $1-\gamma$ proportion of its labels used by $(\alpha+10)n\log{n}$ (Lemma˜9.1). From the previous claim, we know that the coupling does not fail in the 0th phase with probability $1-o(n^{-1})$. Now assume that the coupling first fails in phase $i$. This requires that either $B_{i}>10n\log{n}$ or there is a sufficiently over taxed cut by $B_{i}$. In the former case, since all previous phases did not fail we must have that $10n\log{n}\leq\sum_{j=s}^{s+i}X_{j}\leq\mathbf{X}$. In the latter case, either the former case must be true, or more than a $(1-\beta)$ fraction of edges must have been taken from some cut before $(\alpha+10)n\log{n}$ rounds. Iterating this argument, we find that for any phase to fail, we require that either $\mathbf{X}\geq 10n\log{n}$ or that there exists a set $T$ such that $\sum_{e\in T\times V\setminus T}\frac{|\lambda^{\prime}_{[0,(\alpha+10)n\log{n}]}(e)|}{r|T|(n-|T|)}>\gamma$, the probability of which is at most $o(n^{-1})$ via ˜7, Lemma˜9.1 and a union bound.
@@
As in the equivalent upper bound, our strategy shall be to bound the time between vertices being added to the reachability ball via suitably nice geometric random variables. Similar to before we will define $A_{t}=|\mathcal{B}_{\mathcal{G^{O}}}\left(S,t\right)|$ to be the size of the reachability ball centered at $S$ after $t$ rounds. Further, we will define $B_{i}=\inf\{t>0:A_{t}\geq\frac{n}{2}+i\}$. Our goal is to lower bound $B_{i+1}-B_{i}$ via the independent random variable $Y_{i}\sim\text{Geo}\left(p_{i}\right)$ for $p_{i}=(1+\beta)\frac{2(\frac{n}{2}+i)(\frac{n}{2}-i)}{n(n-1)}$
@@
We construct the following coupling between $B_{i+1}-B_{i}$ and $Y_{i}$, where we denote by $\text{Cross}(i)=\mathcal{B}_{\mathcal{G}^{O}}\left(S,B_{i}\right)\times(V\setminus\mathcal{B}_{\mathcal{G}^{O}}\left(S,B_{i}\right))$ the cut between the reachability ball of $S$ at time $B_{i}$ and the rest of the vertices. First, if $B_{i}>2n\log{n}$ we immediately terminate the coupling and sample trials for $Y_{i}$ independently at random. Otherwise, we define $\text{CrossCount}(i)=\sum_{e\in\text{Cross}(i)}r-|\lambda^{\prime}_{[0,B_{i}]}|$ to be the number of labels left crossing the cut. We reveal each snapshot of $\mathcal{G}^{O}$ in turn, and perform a trial for $Y_{i}$.
@@
By construction we have that either, the coupling fails or the number of trials before a success for $Y_{i}$ ($Y_{i}$) is less than or equal to the number of edges revealed until one crossing the cut $(B_{i+1}-B_{i})$ is. Thus either we have that $Y_{i}<B_{i+1}-B_{i}$ or alternatively $B_{i+1}>2n\log{n}$ (from Lemma˜5.4 probability $o(1)$). Thus, we find that $Y_{i}<B_{i+1}-B_{i}$ with high probability via a union bound over the failure of the coupling.
We will now consider the mass coupling formed by sampling each $Y_{i}$ based on the same sample of $\mathcal{G}$, and show that with sufficiently high probability $Y_{i}<B_{i+1}-B_{i}$ for all $i\in\{0,...,\frac{n-2}{2}\}$. Immediately, by construction we find that each $Y$ random variable is independent of the rest but that have a very high degree of dependence between the failure condition for each individual coupling. For all $i\in\{0,...,\frac{n-2}{2}\}$ we have that $B_{i+1}<2n\log{n}$ if and only if $B_{\frac{n}{2}}<2n\log{n}$. Thus, applying Lemma˜9.3, we have that the probability of ANY of the couplings failing is at most $n^{\frac{\log{2}-1}{5}}$. We therefore have the claim.
@@
$\mathbb{P}\left[\mathbf{Y}\geq\frac{(1-\beta)(1-\delta)n\log{n}}{2(1+\beta)}\right]=1-O(n^{2(1-\beta)(\delta+\log{(1-\delta)})}).$
@@
Proof 10.12.
Proof of Lemma˜9.7 Let $X$ be the number of labels falling before $\frac{\alpha\log{n}}{r(n-1)}$. Each label falls there independently with probability $\frac{\alpha\log{n}}{r(n-1)}$ and so we find that,
$\mathbb{E}[X]=\frac{\alpha}{2}n\log{n}$
Thus from Lemma˜7.3 bound we find that,
Similarly,
In both cases these are $o(n^{-1})$

### v2

@@ 9.2 Lower Bounds
In this section we present a lower bound to match the upper bound of Lemma˜5.5. However, this time, our lives will be significantly easier, as it suffices to consider only the doubling time of sets containing precisely $\frac{n}{2}$ vertices. As before, we need a bound on the doubling time of any specific set, and so make use of the following result analogous to Lemma˜9.3, the proof of which we also defer.
@@
With that established we may now give the main result of the section, a lower bound on the doubling time of $\mathcal{G}^{O}$. See 5.5 The key notion for the proof of this result is the idea of a “blocker vertex”, i.e. a vertex who is not yet reachable by any member of some large set of vertices. By applying Lemma˜9.5 and Markov’s inequality, followed by a simple counting argument, we are able to show that at least a constant fraction of vertices must be blockers around time $\frac{n\log{n}}{2}$. Then, via drift analysis, we are able to show that at least one blocker vertex will not be included in an edge before time $(1-\epsilon)n\log{n}$. Given the fact that the doubling time is lower bounded by the time when the last vertex stops being a blocker, and that a vertex must be included in an edge in order to stop being a blocker, we immediately obtain the bound.
@@
Let $X$ be the number of sets of vertices of size $\frac{n}{2}$ which have a doubling time of at most $\frac{(1-\alpha)n\log{n}}{2(1+\alpha)}$ for some $\alpha>0$ small. We know from Lemma˜9.5 (where the $O(n^{\frac{\log{2}-1}{5}})$ vanishes into the term depending on $\alpha$) and the linearity of expectation that for some constant $c>0$,
@@
Therefore with probability $1-\frac{1}{\log{n}}$ there are at least $(1-cn^{(1-\alpha)(\alpha+\log{(1-\alpha}))})\binom{n}{\frac{n}{2}}$ sets with a doubling time of at least $\frac{(1-\alpha)n\log{n}}{2(1+\alpha)}$.
Now, a set of size $\frac{n}{2}$ has not yet doubled its reachability ball, if and only if, there is some vertex outside of the set that it cannot yet reach. We call such an unreachable vertex a “blocker” vertex for that set. Furthermore, any individual blocker vertex can block at most $\binom{n-1}{\frac{n}{2}}=\frac{1}{n}\binom{n}{\frac{n}{2}}$ sets (precisely those that do not contain it). Since we know that with high probability there are $(1-n^{-\beta})\binom{n}{\frac{n}{2}}$ sets which are blocked for some $\beta>0$, for any constant $0<d<1$ it holds that there must be at least $dn$ blocker vertices.
@@
There exists at least one blocker vertex that does not receive an edge between time $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}$ and $(1-\epsilon)n\log{n}$ with probability 1o(1)1-o(1).
Proof 9.7.
Let $B$ be a set of $dn$ vertices chosen arbitrarily at time $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}$ to either be a subset or super-set of the set of the currently blocking vertices. Let $Y_{t}$ be the set of vertices from $B$ that have not yet received an edge between time $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}$ and $\frac{(1-\alpha)}{2(1+\alpha)}n\log{n}+t$, and $\mathcal{F}_{t}$ be its corresponding filtration. We have the following two bounds, for $t<2n\log{n}$ and any fixed $\gamma>0$ immediately,
@@
The main claim follows immediately from this, the consolidation of constant terms and the observation that if there is at least one blocker vertex that has not received an edge, there must be at least one set that has not yet doubled its size.
@@ 9.3 Tidying Up
In this section, we prove our main theorem on the doubling time in the continuous model. In order to achieve this we will convert our upper and lower bounds on the order discretization of a sample from the continuous model back to the undiscretized version. This is made quite easy by the strong concentration of the number of labels occurring within a given interval. In fact, since $\psi$ is degenerate we obtain the ever pleasant binomial distribution. We are able, therefore, to obtain the following simple precursor, showing that with high probability the number of labels appearing in the initial interval is highly concentrated.
@@
For $\alpha,\beta>0$ with $\beta$ small, the following holds, for any $\mathcal{G}=(V,E,\lambda)$,
$\mathbb{P}\left[\sum_{e\in E}|\lambda^{O}_{[0,\frac{\alpha\log{n}}{r(n-1)}]}(e)|\leq\frac{\alpha(1-\beta)n\log{n}}{2}\right]=o(n^{-1}).$
@@
$\mathbb{P}\left[\sum_{e\in E}|\lambda^{O}_{[0,\frac{\alpha\log{n}}{r(n-1)}]}(e)|\geq\frac{\alpha(1+\beta)n\log{n}}{2}\right]=o(n^{-1}).$
Proof 9.9.
Let $X$ be the number of labels falling before $\frac{\alpha\log{n}}{r(n-1)}$. Since each label falls there independently with probability $\frac{\alpha\log{n}}{r(n-1)}$, we find that,
$\mathbb{E}[X]=\frac{\alpha n\log{n}}{2}$
Thus from a Chernoff bound (Lemma˜7.3) we find that,
Similarly,
In both cases these are $o(n^{-1})$
The proof of the result itself is then simply an application of Lemma˜9.8 to the bounds implied by Lemma˜5.5 and Lemma˜5.4.
Proof 9.10 (Proof of Theorem˜3.5).
@@
Furthermore, we know from Lemma˜9.8 that for any $\delta>0$ the $(1-\epsilon)n\log{n}$th label falls after $\frac{2(1-\delta)(1-\epsilon)\log{n}}{r(n-1)}$ and that the $(1+\epsilon)n\log{n}$th labels falls before $\frac{2(1+\delta)(1+\epsilon)\log{n}}{r(n-1)}$ with probability 1o(1)1-o(1). Thus taking a union bound over the failure probabilities, and $\alpha>\delta+\epsilon+\delta\epsilon$, we obtain the claim.
@@ 10.1.1 First Moment
We first consider the case where $\beta>1$. By linearity, we obtain,
@@
From Lemma˜8.8 we have,
@@
Since $\sum_{\tau\in\mathcal{T}_{S}}\mathbb{P}{[\tau(e)\leq l_{e}]}=\mathbb{E}[\psi]=r$ is both finite and absolutely convergent, by the repeated application of distributivity over absolutely convergent sequences, we obtain,
@@
From Lemma˜8.8 we have,
@@ 10.1.2 Second Moment
Proof 10.2 (Proof of Lemma˜8.22 in the case where $\beta>1$).
@@
Proof 10.3 (Proof of Lemma˜8.22 in the case where $\beta=1$).
@@
$\mathbb{P}{[\mathbf{F}_{(1),T}=0]}\leq\frac{(\frac{\sqrt{2}r}{T})^{2|E_{H}|}n^{2|V_{H}|}(2|V_{H}|)^{4|V_{H}|}\left(1+|E_{H}|(T-1)\right)\sum_{I\subseteq H}n^{-|V_{I}|}\left(\frac{2r}{T}\right)^{-|E_{I}|}}{\left(r^{|E_{H}|}T^{1-|E_{H}|}|V_{H}|!\binom{n}{|V_{H}|}\right)^{2}}$
@@
$O(1)\cdot\frac{1+|E_{H}|(T-1)}{T^{2}}\sum_{I\subseteq H}n^{-|V_{I}|}\left(\frac{2r}{T}\right)^{-|E_{I}|}$
@@ 10.2 Doubling Times
$p=\frac{2}{rn(n-1)-2t}\cdot\left(rs(n-s)-\sum_{e\in\text{Cut}}\lambda^{O}_{[0,t)}(e)\right)$
For $n$ sufficiently large, any fixed $S$, $t\leq\alpha n\log{n}$ and $\gamma>0$, this is trivially upper bounded by,
@@
The result then follows from a simple rearrangement, the fact that $r\geq 1$ and that
@@
With that established, we can start working with the random temporal graph itself. We define $A_{t}=|\mathcal{B}_{\mathcal{G}^{O}_{[\alpha n\log{n},\infty)}}\left(S,\alpha n\log{n}+t\right)|$ to be the size of the reachability ball centered at $S$ after $t$ rounds. Further, we define $B_{i}=\inf\{t\geq 0:A_{t}\geq s+i\}$. We will take a slightly different (but equivalent) view of the process sampling $\mathcal{G}^{O}$: Each edge begins with $r$ tokens. Then on each round, an edge is chosen to become active and discard a token with probability proportional to its remaining tokens, the process concluding once all tokens have been sampled.
@@
We construct the coupling as follows. If there exists a set $T$ such that $\sum_{e\in T\times V\setminus T}\frac{|\lambda^{O}_{[0,\alpha n\log{n}]}(e)|}{r|T|(n-|T|)}>\gamma$ we immediately terminate the coupling, declare it to have failed and sample $X_{s}$ independently of $\mathcal{G}$. If this is not the case, we begin on round $\alpha n\log{n}$ and work round by round. On the $t$th subsequent round,
@@
      $q_{t}=\frac{n(n-1)-2\alpha n\log{n}-2t}{n(n-1)}\cdot\frac{(1-\gamma)rs(n-s)}{rs(n-s)-\sum_{e\in S\times V\setminus S}|\lambda^{O}_{[0,\alpha n\log{n}]}(e)|}$
@@
By construction, we have that, unless the coupling fails, $X_{s}$ samples its first successful trial no earlier than an edge is revealed crossing the cut. However, the coupling only fails if Lemma˜9.1 fails and so we have the claim.
@@
We shall chain together ksk-s variations of the previous coupling and use the good behavior of each one to ensure the good behavior of the next. The coupling is divided into ksk-s phases such that, for $i\in\{0,...,k-s-1\}$, the $i$th phase begins on round $B_{i}$. In the $i$th phase we couple $B_{i+1}-B_{i}$ and $X_{s+i}$, taking
@@
to be the set of edges crossing the cut between the reached and unreached vertices. Firstly, if either of the following hold, we immediately terminate the coupling and sample $X_{s}+i$ independently at random:
@@
    There exists a set $T$ such that $\sum_{e\in T\times V\setminus T}\frac{|\lambda^{O}_{[0,\alpha n\log{n}+B_{i}]}(e)|}{r|T|(n-|T|)}>\gamma$ (i.e. the cut between $T$ and $V\setminus T$ has had too many labels sampled from it).
@@
    $B_{i}>(\alpha+10)n\log{n}$ (i.e. the process has taken improbably long already).
Then, if neither of these occur, we work round by round starting at $\alpha n\log{n}+B_{i}$ and on the $t$th subsequent round.
@@
      $q_{t}=\frac{n(n-1)-2\alpha n\log{n}-2t-2B_{i}}{n(n-1)}\cdot\frac{(1-\gamma)rs(n-s)}{rs(n-s)-\sum_{e\in S\times V\setminus S}|\lambda^{O}_{[0,\alpha n\log{n}+B_{i}]}(e)|}$
@@
  • If an edge has been revealed from $\text{Cross}(i)$, we sample the $t$th trial for $X_{s+i}$ according to its standard success probability, independently of $\mathcal{G}$.
Now note that unless the coupling fails for some $i\in\{0,...,k-s-1\}$ we have that $B_{i+1}-B_{i}\leq X_{s+i}$ for all of the same $i$, and thus $B_{k-s}\leq\mathbf{\bar{X}}$. It therefore remains to show that the coupling does not fail with high probability.
There are two possibilities, either the coupling fails in phase 0 or it fails in a later phase. Since we know from the previous lemma that the coupling succeeds in the zeroth phase with probability 1o(1)1-o(1), any substantial failure probability must occur later. Accordingly, let $j\in\{1,...,k-s-1\}$ be the first phase in which the coupling fails. This requires that either, $B_{j}>10n\log{n}$ or that there is a cut that has been over-sampled by time $\alpha n\log{n}+B_{j}$. In turn, this requires that either $B_{j-1}+X_{s+j-1}>10n\log{n}$ (and therefore $\mathbf{X}>10n\log{n}$) or that there is a cut that has been over-sampled by time $(\alpha+10n\log{n})$. However, from ˜7 and Lemma˜9.1 respectively, we know that both occur with probability at most $o(n^{-1})$. Therefore, we have that in order for the coupling to fail, at least one of three events must occur, each of which does so with probability at most $o(n^{-1})$. It thus follows via a union bound over the failure probabilities, that the coupling succeeds with probability $1-o(n^{-1})$. As such, we have the claim.
@@
As in the equivalent upper bound, our strategy is to bound the time between vertices being added to the reachability ball via suitably nice geometric random variables. Similar to before, we define $A_{t}=|\mathcal{B}_{\mathcal{G^{O}}}\left(S,t\right)|$ to be the size of the reachability ball centered at $S$ after $t$ rounds. Further, we define $B_{i}=\inf\{t>0:A_{t}\geq\frac{n}{2}+i\}$. Our goal is to lower bound $B_{i+1}-B_{i}$ via the independent random variable $Y_{i}\sim\text{Geo}\left(p_{i}\right)$ for $p_{i}=(1+\beta)\frac{2(\frac{n}{2}+i)(\frac{n}{2}-i)}{n(n-1)}$
@@
We construct the following coupling between $B_{i+1}-B_{i}$ and $Y_{i}$, where we denote by $\text{Cross}(i)=\mathcal{B}_{\mathcal{G}^{O}}\left(S,B_{i}\right)\times(V\setminus\mathcal{B}_{\mathcal{G}^{O}}\left(S,B_{i}\right))$ the cut between the reachability ball of $S$ at time $B_{i}$ and the rest of the vertices. Firstly, if $B_{i}>2n\log{n}$ we immediately terminate the coupling and sample trials for $Y_{i}$ independently at random. Otherwise, we define $\text{CrossCount}(i)=\sum_{e\in\text{Cross}(i)}r-|\lambda^{O}_{[0,B_{i}]}|$ to be the number of labels left crossing the cut and reveal each snapshot of $\mathcal{G}^{O}$ in turn. On the $t$th subsequent round,
@@
By construction we have that, either, the coupling fails or the number of trials before a success for $Y_{i}$ is less than or equal to the number of edges revealed until one crossing the cut $(B_{i+1}-B_{i})$ is. Therefore, either we have that $Y_{i}<B_{i+1}-B_{i}$ or, alternatively, $B_{i+1}>2n\log{n}$ (from Lemma˜5.4 probability $o(1)$). Thus, we find that $Y_{i}<B_{i+1}-B_{i}$ with high probability via a union bound over the failure of the coupling.
We will now consider the mass coupling formed by sampling each $Y_{i}$ based on the same sample of $\mathcal{G}$, and show that with sufficiently high probability $Y_{i}<B_{i+1}-B_{i}$ for all $i\in\{0,...,\frac{n-2}{2}\}$. Immediately, by construction we find that each $Y$ random variable is independent of the rest but there is a very high degree of dependence between the failure condition for each individual coupling. In particular, for all $i\in\{0,...,\frac{n-2}{2}\}$, we have that $2n\log{n}<B_{i+1}$ implies that $B_{\frac{n}{2}}>2n\log{n}$. Thus, applying Lemma˜9.3, we have that the probability of ANY of the couplings failing is at most $n^{\frac{\log{2}-1}{5}}$. We therefore have the claim.
@@
$\mathbb{P}\left[\mathbf{Y}\leq\frac{(1-\beta)(1-\delta)n\log{n}}{2(1+\beta)}\right]=1-O(n^{2(1-\beta)(\delta+\log{(1-\delta)})}).$